In a normal distribution 15% of items are below 35 and 10% are above 65. Find the mean.
1
Expert's answer
2020-08-12T17:35:23-0400

Let "X" be a normal random variable with mean "m" and variance "\sigma^2" such that

"P(X<35)=15\%=0.15" and "P(X>65)=10\%=0.1", then

"P(X<65)=1-0.1=0.9."

Since "X" is a normal random variable, then

"Z=\frac{X-m}{\sigma}" is the standard normal random variable.

"P(X<35)=P(\frac{X-m}{\sigma}<\frac{35-m}{\sigma})=\\
P(Z<\frac{35-m}{\sigma})=0.15,"

"P(X<65)=P(\frac{X-m}{\sigma}<\frac{65-m}{\sigma})=\\
P(Z<\frac{65-m}{\sigma})=0.9."

Using quantile function "Q" for the standard normal random variable,

"\frac{35-m}{\sigma}=Q(0.15)=-1.04",

"\frac{65-m}{\sigma}=Q(0.9)=1.28",

hence

"35-m=-1.04\sigma",

"65-m=1.28\sigma",

"65-m-(35-m)=1.28\sigma+1.04\sigma,"

"2.32\sigma=30, \sigma=\frac{30}{2.32}\approx12.931,"

"65-m=1.28\cdot12.931\approx 16.552, \\
m=65-16.552=48.448."

Answer: "48.448."